OpenStudy (anonymous):

if sec t=-5 compute sin(-t)+cos(-t) help!!

OpenStudy (anonymous):

sin(-t)=-sin(t) cos(-t)=cos(t) sin(-t)+cos(-t)=Cos (t) - Sin(t)

OpenStudy (anonymous):

i wrote the problem wrong its if sec x=-5 not t=-5..does this change ur response?

OpenStudy (anonymous):

see... sint=\[\sqrt{1-\cos^2t}\]\[=\sqrt{1-(1/\sec^2t)}\] and sin(-t)=-sint=\[\sqrt{1-(1/25)}\]=\[\sqrt{24}/5\] lly cos(-t)=cost=(-1/5) so.. sin(-t)+cos(-t)=\[\sqrt{24}/5+(-1/5)\]=\[(\sqrt{24}-1)5\] hope its correct

OpenStudy (anonymous):

if you change t to x in sect...then in that case you need to also change t x in cos and sin....in that case it makes no difference..